Ann had $198 more than her sister. After their mother gave Ann $20 and her sister $60, Ann had twice as much money as her sister. How much money did Ann have at first?
Answer by lwsshak3(11628)

Ann had $198 more than her sister. After their mother gave Ann $20 and her sister $60, Ann had twice as much money as her sister. How much money did Ann have at first?
***
let x=money Ann's sister had at first
x+198=money Ann had at first
..
x+198+20=2(x+60)
x+218=2x+120
x=98
x+198=296
How much money did Ann have at first? $296